OpenStudy (anonymous):

How do I find the derivative of this function using the definition of a derivative ? f(x)=4x^2-3x^2+2x-pi Is finding the derivative of a function basically plugging the function into the formula then simplifying it?

OpenStudy (anonymous):

Derivative of a function can be found by using the below formula \[\frac{ d }{ dx}[x^n] = nx^{n-1}\] Example \[\frac{ d }{ dx }[4x^2] = 2*4x^{2-1} = 8x\]

OpenStudy (anonymous):

on our test we cannot use differential formulas we have to solve the derivative using the definition which is f(x)-f(a)/x-a. The only other tool we can use is limit laws

OpenStudy (anonymous):

Okay then you got to use the below definition of derivate \[\frac{dy}{dx}[x]= \lim_{h \rightarrow \infty}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

well yes, I plug it in and I get this far : \[\frac{( 4x^3-3x^2+2x-\pi +h)-4x^3+3x^2+2x-\pi}{ h }\] and I made a typo, the first part is 4x^3 not 4x^2

OpenStudy (anonymous):

I don't know how to simplify is further or how to find the derivative of each individual part

zepdrix (zepdrix):

Woops you didn't plug it in correctly,\[\large\rm f(\color{orangered}{x})=4(\color{orangered}{x})^3-3(\color{orangered}{x})^2+2(\color{orangered}{x})-\pi\]So when evaluated at x+h,\[\large\rm f(\color{orangered}{x+h})=4(\color{orangered}{x+h})^3-3(\color{orangered}{x+h})^2+2(\color{orangered}{x+h})-\pi\]

zepdrix (zepdrix):

So then,\[\large\rm f'(x)=\lim_{h\to0}\frac{f(x+h)-f(x)}{h}\]gives us,\[\rm f'(x)=\lim_{h\to0}\frac{[4(x+h)^3-3(x+h)^2+2(x+h)-\pi]-\left[4x^3-3x^3+2x-\pi\right]}{h}\]

zepdrix (zepdrix):

You'll have to expand out the cubic and square terms, which is going to be a little bit of a pain.

OpenStudy (anonymous):

ok well I multiplied it all out and added/crossed out like terms and I was left with this: \[\frac{ 4h^2+3h^2+2h }{ h}\] then I factored out an h and was left with this unfactorable expression: \[4h^2+3h+2\]

OpenStudy (anonymous):

apparently the answer is supposed to be 12x^2-6x+2

OpenStudy (anonymous):

im not even left with x as a variable im left with h's

zepdrix (zepdrix):

what 0_o

zepdrix (zepdrix):

Like if you look at the first term,\[\large\rm 4(x+h)^3=4[x^3+3x^2h+3xh^2+h^3]\]Certainly that first term, 4x^3 is going to cancel out with something, but not the other stuff. I don't understand how you ended up with no x's >.< hmm

zepdrix (zepdrix):

\[\large\rm -3(x+h)^2=-3[x^2+2xh+h^2]\]

zepdrix (zepdrix):

Lot of stuff to group together D: oh boy

OpenStudy (anonymous):

well I figured that \[-3(x+h)^2 =( -3x^2)+(-3h^2).\]

OpenStudy (anonymous):

I see what you did.....and yeah that is going to take me more time to expand

OpenStudy (anonymous):

I got to : \[\frac{ 12x^2h+12x^2+4h^3-6xh-3h^2 }{ h}\] can I cross out 1 h in every monomial? like this? \[\frac{ 12x^2h+12x^2+4h^3-6xh-3h^2 }{ h}\] --> \[12x^2+12hx+4h^2-6x-3h\]

OpenStudy (usukidoll):

I remember definition of the derivative and it is a royal pain... the h must cancel, so factor all terms with the h and cross it out . then whatever h is remaining take the limit. when h = 0. it's almost there, but it's missing a term the shortcuts (to find the derivative the easy which I don't know why this method isn't taught first) is a great way to check your work f(x)=4x^3-3x^2+2x-pi all constants without a variable will derive to 0. so see that pi over there? the derivative of pi is 0 because we treat pi as a constant f'(x) = 12x^2-6x+2 which is almost correct but somehow the 2 is dropped during the process

OpenStudy (usukidoll):

whoops take the limit when h approaches 0 X>X

OpenStudy (usukidoll):

\[\rm f'(x)=\lim_{h\to0}\frac{[4(x+h)^3-3(x+h)^2+2(x+h)-\pi]-\left[4x^3-3x^2+2x-\pi\right]}{h} \] \[\rm f'(x)=\lim_{h\to0}\frac{[4(x+h)^3-3(x+h)^2+2(x+h)-\pi]+ \left[-4x^3+3x^2-2x+\pi\right]}{h} \] \[\rm f'(x)=\lim_{h\to0}\frac{[4x^3+12x^2h+12hx^2+4h^3-3x^2-6xh-3h^2+2x+2h-\pi]+ \left[-4x^3+3x^2-2x+\pi\right]}{h} \]

OpenStudy (usukidoll):

4x^3, 3x^2,2x, and pi should all cancel

OpenStudy (usukidoll):

\rm f'(x)=\lim_{h\to0}\frac{[12x^2h+12hx^2+4h^3-6xh-3h^2+2h-\pi]}{h}

OpenStudy (usukidoll):

\[\frac{12x^2h+12h^2x+4h^3-6xh-2h^2+2h}{h}\] factor all the h's \[\frac{(h)[12x^2+12hx+4h^2-6x-2h+2]}{h}\] now the h does cancel and you're left with \[12x^2+12hx+4h^2-6x-2h+2 \] take the limit as h approaches 0 for any term with an h left over. those terms should disappear and you should have the derivative. I know I made an error and didn't realize it until the end but I can't edit that. Also, my latex broke and my batteries for my mouse dies.

OpenStudy (usukidoll):

*died rather.. but on one of the 2's there's an h which goes away when you take the limit but then there's the other 2 so that saves the day

OpenStudy (anonymous):

thanks a lot ! I have another quick question. I need to find the derivative for this using the definition : \[\frac{ x^2 }{ 3}-\frac{ 3 }{ x^2 }\] When I plug it into the definition is should look like this right? : \[\frac{ \frac{ (x+h)^2 }{ 3 }-\frac{ 3 }{( x+h)^2} }{ h }\]

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

I am away from my desktop and need to look at this later. I think it would help if you combine the fraction first and then apply the definition of the derivative

OpenStudy (usukidoll):

Or maybe not. There would've been x^4 involved and that would involve the binomial theorem for the x+h parts

OpenStudy (usukidoll):

using the definition of the derivative \[\frac{f(x+h)-f(x)}{h}\] we should have \[\huge\frac{\frac{ (x+h)^2 }{ 3}-\frac{ 3 }{ (x+h)^2 }-(\frac{ x^2 }{ 3}-\frac{ 3 }{ x^2 })}{h}\]

OpenStudy (usukidoll):

so the gcf is\[ (x+h)^2 (x^2)(3) \] for the upper fraction... wow this is crazy with the definition version.

OpenStudy (usukidoll):

sigh all that work for this \[\LARGE f(x) = \frac{1}{3}x^2-3x^{-2}\] \[\LARGE f'(x) = \frac{2}{3}x-(-2)3x^{-2-1}\] \[\LARGE f'(x) = \frac{2}{3}x+6x^{-3}\] \[\LARGE f'(x) = \frac{2}{3}x+\frac{6}{x^3}\]